f two six-sided dice are rolled, what is the probability that the total number of spots showing is 4?

a.

1/6

b.

1/18

c.

1/12

d.

1/3

e.

1/36

-------------------------------------------------------------------------------------------------------------------------------------------------------

Red-Green colorblindness is a X-linked recessive disorder. Huntington’s is an autosomal dominant disorder. Susan’s father is color blind and has Huntington’s disease and is heterozygous for the Huntington’s allele.   Her mother is normal and has no history of colorblindness or Huntington’s in her family. Her brother has developed Huntington’s but is not colorblind. Her husband Greg is healthy and has no incidence of either disorder in his family.   Susan is not colorblind and has not yet be tested to see if she has inherited the Huntington’s allele. What is the chance that a child of Susan and Greg will be a girl with normal vision but having Huntington’s?

a.

1/16

b.

1/8

c.

1/64

d.

1/4

e.

1/32
Answer #1

1. c. 1/12

n(s) = 36

The probability of getting total spot showing 4 is n(A) = 3 {(1,3), (2,2), (3,1)}

Therefore, n(A) / n(S) = 3/36 = 1/12

2. d. 1/4

When a colour blind man marries with a normal visioned woman, they will produce normal visioned daughters and sons in F1. The marriage between F1 normal visioned woman and normal visioned male, will produce in F2 two normal visioned females, one normal visioned male and one colourblind male.